# Identifying Network Structure of Nonlinear Dynamical Systems: Contraction and Kuramoto Oscillators

### v1

@@ Abstract
In this work, we study the identifiability of network topologies for networked nonlinear systems when partial measurements of the nodes are taken. We explore scenarios where different candidate topologies can yield similar measurements, thus limiting identifiability. To do so, we apply the contraction theory framework to facilitate comparisons between candidate topologies. We show that semicontraction in the observable space is a sufficient condition for two systems to become indistinguishable from one another based on partial measurements. We apply this framework to study networks of Kuramoto oscillators, and discuss scenarios in which different topologies (both connected and disconnected) become indistinguishable.
@@ I Introduction and Motivation
Many structures in the natural world (e.g., brains, social networks) can be described as networks of interconnected dynamical systems [1, 2]. Thus, there is significant interest in the identification of network topologies from time-series measurements of nodes in these networks [3, 4, 5]. In the context of neuroscience, estimating the topology of neuronal circuits and brain networks from partial measurements of the nodes is a longstanding problem [6, 2]. Moreover, given a set of measurements, different methods may identify different network topologies [3, 7]. In networked nonlinear systems, behavior such as synchrony [8, 9] can cause measurements from several different systems to become identical, presenting challenges for identification [5] (see Fig. 1). Partial measurements of the nodal states can lead to further challenges in identifying the true underlying network topology.
@@
In this paper, we leverage contraction theory to compare the dynamics of networked nonlinear systems [^1] For results on network topology identification in linear setting, see companion paper [7]. with different topologies. Contraction theory provides a framework to study the dynamics of nonlinear systems with varying initial conditions, and has been used to discuss the behavior of interconnected systems and networks of oscillators [10, 11, 12]. We are specifically interested in the limits of topology identification in the nonlinear setting, that is: which network topologies are impossible to distinguish based on partial measurement? In the context of contraction, this translates to: which network topologies exhibit contraction in the observable space? This diverges from standard notions of contraction, which assume full observation and focus on one system with varying initial conditions. In contrast, our work focuses on contraction in the observable space, i.e., different systems exhibiting similar observed dynamics (we provide a rigorous definition in Section II).
@@
Figure 1: Various network topologies can have observed trajectories that contract to one another making them indistinguishable from each other. We derive sufficient conditions on the dynamics for this phenomenon to occur, and in doing so, we are able to generate other candidate topologies that behave similarly (i.e., have the same observed trajectories).
@@
We begin by developing a framework to analyze the space of network topologies that are indistinguishable based on partial measurements of the nodes’ states, and establish sufficient conditions for indistinguishability (Section II). We then show how this framework can be applied to nonlinear (Section II-A) and linear (Section II-B) networked systems. We then apply this theory to Kuramoto oscillator networks, and show how one can generate a variety of candidate Kuramoto oscillator networks that are indistinguishable based on partial measurements (Section III). Concluding remarks are provided in Section IV.
@@
Notation.  We denote matrices ($A$) and vectors ($x$) with uppercase and lower case letters respectively. We denote the vector of all 1’s with $\mathbbm{1}$. The norm with respect to a matrix $M$ is denoted by $\|x\|_{M}=\sqrt{x^{\top}Mx}$, and the $\ell_{\infty}$ norm as $\|x\|_{\infty}=\max_{i}|x_{i}|$. The partial derivative of a function $f$ with respect to a variable $s$ is denoted by $\partial_{s}f$ and the time derivative by $\frac{df}{dt}$ or equivalently $\dot{f}$. The pseudoinverse of a matrix $A\in\mathbb{R}^{m\times n}$ with $\text{rank}(A)=m$ is given as $A^{\dagger}=A^{\top}(AA^{\top})^{-1}$. The spectral abscissa of a matrix is given as $\alpha(A)=\max_{i}\{\text{Re}(\lambda_{i}(A))\}$. Lastly, $\mathcal{R}(\cdot)$ and $\mathcal{N}(\cdot)$ correspond to the range and nullspace of their respective arguments, and $\oplus$ is the direct sum of two vector spaces.
@@ II Contraction Theory for System Comparisons
We consider two systems, $\Sigma$ and $\tilde{\Sigma}$, evolving under the following dynamics
$\Sigma:\begin{cases}\dot{x}=f(x)&x(0)=x_{0}\\ y=Cx\end{cases},$ (1)
$\tilde{\Sigma}:\begin{cases}\dot{\tilde{x}}=g(\tilde{x})&\tilde{x}(0)=\tilde{x}_{0}\\ \tilde{y}=C\tilde{x}\end{cases}$. (2)
Here $x,\tilde{x}\in\mathbb{R}^{n}$ encode the states of each of the systems and, $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ are nonlinear functions encoding the dynamics. Lastly, $y,\tilde{y}\in\mathbb{R}^{p}$ are measurements of the systems through the observation matrix $C\in\mathbb{R}^{p\times n}$. We note that the choice to make linear measurements of the states is analogous to neuroscience settings where microelectrode arrays are used to capture single neuron or multiple neuron recordings in a superposition. We now formalize the notion of indistinguishability.
@@
To compare the trajectories of the systems (1) and (2), and determine whether they become indistinguishable, we consider the auxiliary system inspired by [13]:
$\Sigma_{\text{aux}}:\begin{cases}\dot{\mathbf{x}}=f(\mathbf{x})+s\Delta(\mathbf{x})&\mathbf{x}(0)=(1-s)x_{0}+s\tilde{x}_{0}\\ \mathbf{y}=C\mathbf{x}\\ \Delta(\mathbf{x})=g(\mathbf{x})-f(\mathbf{x})\end{cases}.$ (3)
@@
The auxiliary system defined in (3) is a convex combination of the dynamics in (1) and (2). In [10, Example 2.2] it is shown that a convex combination of contracting systems tend to a common trajectory or equilibrium point shared among them.
We now measure the time-varying length [13, 14] with respect to some positive definite matrix $M(x)\in\mathbb{R}^{p\times p}$ between the trajectories of (1) and (2) defined as
$L(t)\coloneq\int_{0}^{1}\|\partial_{s}\mathbf{y}(t,s)\|_{M}ds.$ (4)
With this, we can now study the dynamics of this length that measures a parametrized arc that connects trajectories of the systems (1) and (2). We define the matrix $M_{C}\coloneq C^{\top}MC$.
@@
Let $\mu\in\mathbb{R}_{>0}$ and assume the auxiliary system (3) is contracting with respect to $M_{C}$ with rate $\mu$, namely it satisfies the linear matrix inequality (LMI)
$\frac{\partial f}{\partial\mathbf{x}}^{\top}M_{C}+\dot{M}_{C}+M_{C}\frac{\partial f}{\partial\mathbf{x}}\\ +s\bigg{(}\frac{\partial\Delta}{\partial\mathbf{x}}^{\top}M_{C}+M_{C}\frac{\partial\Delta}{\partial\mathbf{x}}\bigg{)}\preceq-2\mu M_{C},$ (5)
then the time varying length $L(t)$ satisfies the differential inequality
@@
$\mathbf{y}(t,s)=C\bigg{(}\mathbf{x}(0,s)+\int_{\tau=0}^{t}f(\mathbf{x}(\tau,s))+s\Delta(\mathbf{x}(\tau,s))d\tau\bigg{)}.$ (7)
Furthermore, differentiating with respect to $s$ yields
$\partial_{s}\mathbf{y}(t,s)=\\ C\bigg{(}\partial_{s}\mathbf{x}(0,s)+\int_{\tau=0}^{t}\partial_{s}\{f(\mathbf{x}(\tau,s))+s\Delta(\mathbf{x}(\tau,s))\}d\tau\bigg{)}.$ (8)
@@
$\frac{d}{dt}\partial_{s}\mathbf{y}(t,s)=C\frac{\partial f}{\partial\mathbf{x}}\partial_{s}x+C\Delta+s\frac{\partial\Delta}{\partial\mathbf{x}}\partial_{s}x.$ (9)
@@
$\frac{d}{dt}\|\partial_{s}\mathbf{y}\|_{M}^{2}=\partial_{s}\mathbf{x}^{\top}\bigg{\{}\frac{\partial f}{\partial\mathbf{x}}^{\top}M_{C}+\dot{M}_{C}+M_{C}\frac{\partial f}{\partial\mathbf{x}}\\ +s\bigg{(}\frac{\partial\Delta}{\partial\mathbf{x}}^{\top}M_{C}+M_{C}\frac{\partial\Delta}{\partial\mathbf{x}}\bigg{)}\bigg{\}}\partial_{s}\mathbf{x}+2\partial_{s}\mathbf{x}^{\top}M_{C}\Delta.$ (10)
@@
$\frac{d}{dt}\|\partial_{s}\mathbf{y}\|_{M}^{2}\leq-2\mu\|\partial_{s}\mathbf{y}\|_{M}^{2}+2\|\partial_{s}\mathbf{y}\|_{M}\|C\Delta\|_{M}.$ (11)
Simply differentiating (4) and substituting (11) will yield the result (see [14, Theorem 2.3] for details). ∎
@@
If we find a semidefinite matrix, $M\succeq 0$, satisfying (5) we can guarantee contraction in all space except for $\mathcal{N}(M)$. See [12] for discussion on the notion of semicontraction.
Lemma 1 establishes that $L(t)$ decreases and the length remains bounded provided $\Delta(\mathbf{x})$ is bounded $\forall t,s$. We note that this result is analogous to the discussion on disturbed flows in [11].
@@
We note that Lemma 1 is an extension of traditional contraction theory results. Namely, this framework shows the impact of partial observation via the matrix $M_{C}$, and provides an explicit approach on how one can compare the dynamics of different systems.
@@
If the difference between the dynamics, $\Delta(\mathbf{x})$, satisfies $C\Delta(\mathbf{x})=0$ the two systems’ measurements contract to each other with rate $\mu$.
@@
Corollary 2 establishes that if the difference between the dynamics, $\Delta(\mathbf{x})$, is in the nullspace of the observation matrix, $C$, then as $t\rightarrow\infty$ the observations $y(t)$ and $\tilde{y}(t)$ will become indistinguishable. Naturally, it begs the question, what conditions need to be enforced on the Jacobians of $f$ and $\Delta$ in order for the LMI in (5) to be satisfied and hence for the systems in (1) and (2) to become indistinguishable? In order to understand this question, we consider constant matrices $M(x)=M$ and establish the following definition.
@@
We now show that contraction in the observable space is directly related to the indistinguishability of topologies by showing that it is a requirement for (5) to be satisfied. In the context of the auxiliary system, we would require
@@
We begin with the forward implication. Suppose that $\alpha(CAC^{\dagger})\leq-\mu$, then we know that $\exists M\succ 0$ such that the following LMI holds
@@
Pre and post multiplying by $CC^{\top}$ we arrive at
@@
Identifying that $\mathbb{R}^{n}=\mathcal{R}(C^{\top})\oplus\mathcal{N}(C)$ we can decompose any vector $v$ as $v=C^{\top}m+q$ where $q\in\mathcal{N}(C)$. Thus we have that
@@
We now proceed to prove the reverse implication by assuming that (16) holds. Pre and post multiplying (16) by $C^{\top}C\succeq 0$ we arrive at
@@
Identifying that $\mathbb{R}^{p}=\mathcal{R}(C)\oplus\mathcal{N}(C^{\top})$ we can express any vector $v$ as $v=Cm+q$ where $q\in\mathcal{N}(C^{\top})$. Applying a similar argument as made previously
@@
which leads to
$CA^{\top}M_{C}C^{\top}+CM_{C}AC^{\top}+2\mu CM_{C}C^{\top}\preceq 0,$ (23)
furthermore, pre and post multiplying by $(CC^{\top})^{-1}$ establishes
@@
The assumption $\mathcal{N}(C)\subseteq\mathcal{N}(CA)$ is not required for the reverse implication in Theorem 1. Moreover, this assumption amounts to requiring $\mathcal{N}(C)$ to be $A$-invariant. In practice this is not regularly satisfied, but in certain cases (see Section III) can be. In the case $C=I$ it is trivially satisfied and (16) reduces to a Lyapunov equation, and (14) is simply asking for $A$ to be Hurwitz.
Theorem 1 implies that the requirement $\alpha(CAC^{\dagger})<0$ is essential in order for the matrix $M$ to exist. This underscores that the auxiliary system in (3) must be contracting in the observable space for the two systems being compared to have trajectories that contract to each other and hence be indistinguishable. As an immediate requirement, it is necessary to identify whether (15) is satisfied. Notably, the condition given in (15) can be challenging to verify if the Jacobians depend explicitly on $\mathbf{x}$. As a summary, it is necessary to execute the following steps to determine whether two systems are indistinguishable:
@@
    Verify (15) and find a suitable $M\succeq 0$ that satisfies (5) by applying Theorem 1. One may allow for semidefinite matrices for further generality.
@@
    To satisfy Corollary 2 one must determine whether $C\Delta(\mathbf{x})=0$.
II-A Application I: Nonlinear Networks
The results described previously provide a general framework for comparing the observed dynamics of nonlinear systems. We now particularize these results to nonlinear network systems in order to describe how a variety of network topologies can yield dynamics that are similar. Specifically, consider the dynamics of a nonlinear network given as
$\Sigma_{\text{net}}:\begin{cases}\dot{x}=h(x)+W\Phi(x)&x(0)=x_{0}\\ y=Cx\end{cases}.$ (25)
Here $h(x)$ is a vector encoding the internal dynamics of the nodes in the network, specifically, $h(x)$ does not encode any interaction between nodes and all interactions are captured in the coupling term $W\Phi(x)$, where $W$ is a coupling matrix, and $\Phi(x)$ is some nonlinear function applied element-wise to the nodal states. Notably, applying our framework, we can determine perturbations to the network coupling matrix $W+\Delta$ that will yield identical or similar dynamics. The corresponding perturbed system dynamics would be
$\tilde{\Sigma}_{\text{net}}:\begin{cases}\dot{x}=h(x)+(W+\Delta)\Phi(x)&x(0)=x_{0}\\ y=Cx\end{cases}.$ (26)
Moreover, the condition outlined in (15) amounts to verifying whether

### v2

@@ Abstract
In this work, we study the identifiability of network structures (i.e., topologies) for networked nonlinear systems when partial measurements of the nodal dynamics are taken. We explore scenarios where different candidate structures can yield similar measurements, thus limiting identifiability. To do so, we apply the contraction theory framework to facilitate comparisons between different networks. We show that semicontraction in the observable space is a sufficient condition for two systems to become indistinguishable from one another based on partial measurements. We apply this framework to study networks of Kuramoto oscillators, and discuss scenarios in which different network structures (both connected and disconnected) become indistinguishable.
@@ I Introduction and Motivation
Many aspects of the natural world (e.g., brains, social networks) can be described as networks of interconnected dynamical systems [1, 12]. Thus, there is significant interest in the identification of network structures (i.e., topology) from time series measurements of nodal dynamics [14, 7, 3]. In the context of neuroscience, estimating the structure of neuronal circuits and brain networks from partial measurements is a longstanding problem [10, 12]. Moreover, given a set of measurements, different methods may identify different networks [14, 6]. In networked nonlinear systems, behavior such as synchrony [5, 8] can cause measurements from several different systems to become identical, presenting challenges for identification [3] (see Fig. 1). Access to partial measurements of the nodal dynamics exacerbate this challenge.
@@
In this paper, we leverage contraction theory to compare the dynamics of networked nonlinear systems [^1] For results on network structure identification in the linear setting, see companion paper [6]. with different network structures. Contraction theory provides a framework to study the dynamics of nonlinear systems with varying initial conditions, and has been used to discuss the behavior of interconnected systems and networks of oscillators [16, 11, 2]. We are specifically interested in the limits of topology identification in the nonlinear setting, that is: which network structures are impossible to distinguish based on partial measurement? In the context of contraction, this translates to: which network structures exhibit contraction in the observable space? This diverges from standard notions of contraction, which assume full observation and focus on one system with varying initial conditions. In contrast, our work focuses on contraction in the observable space, i.e., different systems exhibiting similar observed dynamics (we provide a rigorous definition in Section II).
@@
Figure 1: Various network structures can have observed trajectories that contract to one another making them indistinguishable from each other. We derive sufficient conditions on the dynamics for this phenomenon to occur, and in doing so, we are able to generate other candidate structures that behave similarly (i.e., have the same observed trajectories).
@@
We begin by developing a framework to analyze the space of networks that are indistinguishable based on partial measurements of the nodes’ states, and establish sufficient conditions for indistinguishability (Section II). We then show how this framework can be applied to nonlinear (Section II-A) and linear (Section II-B) networked systems. We apply this theory to Kuramoto oscillator networks, and show how one can generate a variety of candidate Kuramoto oscillator networks that are indistinguishable based on partial measurements (Section III, IV). Concluding remarks are provided in Section V.
@@
Notation.  We denote the vector of all 1’s with $\mathds{1}$. The norm with respect to a matrix $M$ is $\|x\|_{M}=\sqrt{x^{\top}Mx}$, and the $\ell_{\infty}$ norm is $\|x\|_{\infty}=\max_{i}|x_{i}|$. The partial derivative of a function $f$ with respect to a variable $s$ is denoted by $\partial_{s}f$ and the time derivative by $\frac{df}{dt}$ or equivalently $\dot{f}$. The pseudoinverse of a matrix $A\in\mathbb{R}^{m\times n}$ with $\text{rank}(A)=m$ is $A^{\dagger}=A^{\top}(AA^{\top})^{-1}$. The spectral abscissa of a matrix is given as $\alpha(A)=\max_{i}\{\text{Re}(\lambda_{i}(A))\}$. Lastly, $\mathcal{R}(\cdot)$ and $\mathcal{N}(\cdot)$ correspond to the range and nullspace of their respective arguments, and $\oplus$ is the direct sum of two vector spaces.
@@ II Contraction Theory for System Comparisons
We consider two systems $\Sigma$ and $\tilde{\Sigma}$ evolving under the following dynamics
$\Sigma:\begin{cases}\dot{x}=f(x),&x(0)=x_{0}\\ y=Cx,\end{cases}$ (1)
[…]
Here $x,\tilde{x}\in\mathbb{R}^{n}$ encode the states of each of the systems and, $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ are nonlinear functions encoding the dynamics. Lastly, $y,\tilde{y}\in\mathbb{R}^{p}$ are measurements of the systems through the observation matrix $C\in\mathbb{R}^{p\times n}$. The choice to analyze linear measurements of the states is reasonable as, for instance, in neuroscience settings microelectrode arrays are used to capture single neuron or multiple neuron recordings in a linear superposition. We now formalize the notion of indistinguishability.
@@
To compare the trajectories of the systems (1) and (2) and determine whether they become indistinguishable, we consider the auxiliary system inspired by [9]:
$\Sigma_{\text{aux}}:\begin{cases}\dot{\mathbf{x}}=f(\mathbf{x})+s\Delta(\mathbf{x}),&\mathbf{x}(0)=(1-s)x_{0}+s\tilde{x}_{0}\\ \mathbf{y}=C\mathbf{x},\\ \Delta(\mathbf{x})=g(\mathbf{x})-f(\mathbf{x}).\end{cases}$ (3)
@@
The auxiliary system defined in (3) is a convex combination of the dynamics in (1) and (2). In [16, Ex. 2.2] a convex combination of contracting systems is shown to tend to a shared trajectory or equilibrium point.
We now measure the time-varying length [9, 13] between the trajectories of (1) and (2) with respect to some positive definite matrix $M(\mathbf{x})\in\mathbb{R}^{p\times p}$ defined as
$L(t)=\int_{0}^{1}\|\partial_{s}\mathbf{y}(t,s)\|_{M}ds.$ (4)
We now study the dynamics of this length. For brevity, we define the matrix $M_{C}=C^{\top}MC$.
@@
Assume the auxiliary system (3) is contracting with respect to $M_{C}$ with rate $\mu\in\mathbb{R}_{>0}$, i.e., it satisfies the linear matrix inequality (LMI)
$\big(\tfrac{\partial f}{\partial\mathbf{x}}+s\tfrac{\partial\Delta}{\partial\mathbf{x}}\big)^{\top}M_{C}+\dot{M}_{C}+M_{C}\big(\tfrac{\partial f}{\partial\mathbf{x}}+s\tfrac{\partial\Delta}{\partial\mathbf{x}}\big)\preceq-2\mu M_{C},$ (5)
then the time-varying length $L(t)$ satisfies the differential inequality
@@
$\mathbf{y}(t,s)=C\Big(\mathbf{x}(0,s)+\int_{\tau=0}^{t}f(\mathbf{x}(\tau,s))+s\Delta(\mathbf{x}(\tau,s))d\tau\Big).$ (7)
Differentiating with respect to $s$ yields
$\partial_{s}\mathbf{y}(t,s)=\\ C\Big(\partial_{s}\mathbf{x}(0,s)+\int_{\tau=0}^{t}\partial_{s}\{f(\mathbf{x}(\tau,s))+s\Delta(\mathbf{x}(\tau,s))\}d\tau\Big).$ (8)
@@
$\tfrac{d}{dt}\partial_{s}\mathbf{y}(t,s)=C\tfrac{\partial f}{\partial\mathbf{x}}\partial_{s}\mathbf{x}+C\Delta+sC\tfrac{\partial\Delta}{\partial\mathbf{x}}\partial_{s}\mathbf{x}.$ (9)
@@
$\tfrac{d}{dt}\|\partial_{s}\mathbf{y}\|_{M}^{2}=\partial_{s}\mathbf{x}^{\top}\bigg\{\tfrac{\partial f}{\partial\mathbf{x}}^{\top}M_{C}+\dot{M}_{C}+M_{C}\tfrac{\partial f}{\partial\mathbf{x}}\\ +s\Big(\tfrac{\partial\Delta}{\partial\mathbf{x}}^{\top}M_{C}+M_{C}\tfrac{\partial\Delta}{\partial\mathbf{x}}\Big)\bigg\}\partial_{s}\mathbf{x}+2\partial_{s}\mathbf{x}^{\top}M_{C}\Delta.$ (10)
@@
$\tfrac{d}{dt}\|\partial_{s}\mathbf{y}\|_{M}^{2}\leq-2\mu\|\partial_{s}\mathbf{y}\|_{M}^{2}+2\|\partial_{s}\mathbf{y}\|_{M}\|C\Delta\|_{M}.$ (11)
Differentiating (4) and substituting (11) will lead to the result (see [13, Thm. 2.3] for details). ∎
@@
If we find a semidefinite matrix $M\succeq 0$ satisfying (5) we can guarantee contraction in all space except for $\mathcal{N}(M)$. See [2] for discussion on the notion of semicontraction.
Lemma 1 establishes that $L(t)$ decreases and remains bounded provided $\Delta(\mathbf{x})$ is bounded $\forall t,s$. This result is analogous to the discussion on disturbed flows in [11].
@@
Lemma 1 is an extension of traditional contraction theory results. This framework shows the impact of partial observation via $M_{C}$, and provides an explicit approach on how to compare the dynamics across systems.
@@
If the difference between the dynamics $\Delta(\mathbf{x})$ satisfies $C\Delta(\mathbf{x})=0$ the two systems’ measurements contract to each other with rate $\mu$.
@@
Corollary 2 establishes that if the difference between the dynamics $\Delta(\mathbf{x})$ is in the nullspace of the observation matrix $C$ then as $t\rightarrow\infty$ the observations $y(t)$ and $\tilde{y}(t)$ will become indistinguishable. This motivates determining the conditions that need to be enforced on the Jacobians of $f$ and $\Delta$ in order for the LMI in (5) to be satisfied; which would imply that the systems in (1) and (2) would become indistinguishable. To do so, we consider constant matrices $M(x)=M$ and establish the following definition.
@@
We now show that contraction in the observable space is directly related to the indistinguishability of network structures by showing that it is a requirement for (5) to be satisfied. In the context of the auxiliary system, we require
@@
We begin with the forward implication. Suppose that $\alpha(CAC^{\dagger})\leq-\mu$, then $\exists M\succ 0$ such that the following LMI holds
@@
Pre- and post-multiplying by $CC^{\top}$ we arrive at
@@
Identifying that $\mathbb{R}^{n}=\mathcal{R}(C^{\top})\oplus\mathcal{N}(C)$ we can decompose any vector $v$ as $v=C^{\top}m+q$ where $q\in\mathcal{N}(C)$. Hence,
@@
We now prove the reverse implication by assuming that (16) holds. Pre- and post-multiplying (16) by $C^{\top}C\succeq 0$ we arrive at
@@
Since $\mathbb{R}^{p}=\mathcal{R}(C)\oplus\mathcal{N}(C^{\top})$ we can express any vector $v$ as $v=Cm+q$ where $q\in\mathcal{N}(C^{\top})$. Applying a similar argument as made previously
@@
which implies
$CA^{\top}M_{C}C^{\top}+CM_{C}AC^{\top}+2\mu CM_{C}C^{\top}\preceq 0.$ (23)
Now, pre- and post-multiplying by $(CC^{\top})^{-1}$ establishes
@@
The assumption $\mathcal{N}(C)\subseteq\mathcal{N}(CA)$ is not required for the reverse implication in Theorem 1. Moreover, this assumption amounts to requiring $\mathcal{N}(C)$ to be $A$-invariant. In practice this is not regularly satisfied, but in certain cases (see Section III) can be. When $C=I$ it is trivially satisfied, (16) reduces to a Lyapunov equation and (14) implies $A$ must be Hurwitz.
Theorem 1 implies that satisfying $\alpha(CAC^{\dagger})<0$ is essential in order for the matrix $M$ to exist. This underscores that the auxiliary system in (3) must be contracting in the observable space for the two systems being compared to have trajectories that contract to each other and become indistinguishable. That is, it is necessary to identify whether (15) is satisfied which may be challenging if the Jacobians depend explicitly on $\mathbf{x}$.
As a summary, the following steps will determine whether two systems are indistinguishable:
@@
    Verify (15) and find a suitable $M\succeq 0$ that satisfies (5) by applying Theorem 1. Semidefinite matrices allow for further generality.
@@
    To satisfy Corollary 2, determine whether $C\Delta(\mathbf{x})=0$.
II-A Application I: Nonlinear Networked Systems
The results described previously provide a general framework for comparing the observed dynamics of nonlinear systems. We now particularize these results to nonlinear networked systems to describe how a variety of network structures can yield dynamics that are similar. Specifically, consider the dynamics of a network given as
$\Sigma_{\text{net}}:\begin{cases}\dot{x}=h(x)+W\Phi(x),&x(0)=x_{0}\\ y=Cx.\end{cases}$ (25)
Here $h(x)$ is a vector encoding the self dynamics of individual nodes in the network, i.e., it does not encode any interaction between nodes. Interactions between nodes are captured in the coupling term $W\Phi(x)$, where $W$ encodes the structure of the network, and $\Phi(x)$ is a nonlinear function applied element-wise to the nodal states. Applying our framework, we can determine whether a perturbed network $W+\Delta$ will yield identical or similar dynamics. The corresponding perturbed system dynamics would be
$\tilde{\Sigma}_{\text{net}}:\begin{cases}\dot{x}=h(x)+(W+\Delta)\Phi(x),&x(0)=x_{0}\\ y=Cx.\end{cases}$ (26)
The condition outlined in (15) amounts to verifying whether